OpenStudy (anonymous):

Identify the vertex, axis of symmetry, and min/max value of each. f(x) = -2x^2 - 36x -164 f(x) = -7x^2 + 14x - 15 f(x) = -6x^2 + 12x - 4 f(x) = 10x^2 - 100x +240

OpenStudy (anonymous):

each of these works the same way. find \[-\frac{b}{2a}\] and that will be the first coordinate of the vertex. second is what you get by substitution

OpenStudy (anonymous):

\[f(x) = -2x^2 - 36x -164 \] \[-\frac{b}{2a}=-\frac{36}{4}=-9\] \[f(9)=-2\] range is \[(-\infty, -2]\]

OpenStudy (anonymous):

\[f(x) = -7x^2 + 14x - 15\] \[-\frac{b}{2a}=\frac{14}{2\times 7}=1\] \[f(1)=-8\] vertex is \[(1,-8)\] range is \[(-\infty, -8]\]

OpenStudy (anonymous):

What do you mean by 'range'?